# Dissipating quartets of excitations in a superconducting circuit

### v1

@@ IV Conclusion and outlook
We demonstrate the engineering of a four-photon dissipation channel whose rate far exceeds the intrinsic single-photon loss rate of the target oscillator. Our approach relies on a near Kerr-free Josephson mixing element embedded in a high-impedance environment, mediating genuine high-order interactions while inducing only residual non-linearity on the involved modes. Even though this non-linearity is larger than the engineered dissipation rate in our experiment—preventing us from dissipating quartets of photons from more than a specific oscillator state—we show that it dominantly originates from a renormalization of the SQUID inductive potential by stray inductances in its tracks. We thus estimate that simple design changes would decrease the circuit non-linearity down to a level where it can be dynamically canceled [47]. The four-photon dissipation channel would then become state-agnostic, enabling the dissipative stabilization of a four-legged cat manifold.
@@ Acknowledgments
The authors thank S. Jezouin, R. Lescanne, A. Murani, and W.C. Smith for fruitful discussions on circuit design, J. Palomo, A. Pierret, M. Rosticher and J.-L. Smirr for assistance in microfabrication, and M. Mirrahimi, P. Rouchon and L.-A. Sellem for valuable discussions on dissipation engineering. Our circuit was fabricated within the consortium Salle Blanche Paris Centre and at the Collège de France. This work has been funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101042304), by the French ANR-22-PETQ-0003 grant and by the ANR-22-PETQ-0006 grant under the ‘France 2030’ plan.
@@
with $\gamma_{4}=4g_{4}^{2}/\kappa_{b}$. The coherent states spanning the cat manifold are attractors of this dynamics. Note that this property remains true when the above hierarchy between $\kappa_{b}$ and $g_{4}$ is not respected [27]. In (S1), the term $\alpha^{4}{\bf b^{\dagger}}+\mathrm{h.c.}$ is trivially implemented with a resonant drive on the buffer, so that we now focus on the engineering of the more challenging four-to-one photon exchange interaction.
In practice, we engineer this interaction using a pumped ATS participating in the memory and buffer (see Fig. 1). After Foster decomposition of the linear circuit connected to the ATS [48, 21], we reach the Hamiltonian (2) that we recall here:
@@
If the system is pumped at the frequency $\epsilon_{p}(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$ and after selecting resonant terms, we get in the RWA the static Hamiltonian
$\tilde{{\bf H}}_{4}=g_{4}:P_{a}({\bf a^{\dagger}}{\bf a})P_{b}({\bf b^{\dagger}}{\bf b})~{}{\bf a}^{4}{\bf b^{\dagger}}:~{}+~{}\mathrm{h.c.}$ (S5)
@@
$\begin{split}P_{a}(X)=\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{a}^{ZPF}}^{2n}}{n!(n+4)!}X^{n}\\ P_{b}(X)=\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{b}^{ZPF}}^{2n}}{n!(n+1)!}X^{n}\end{split}$ (S6)
@@
Figure S4: Sources of non-linearity. a) Ideal circuit (no source of non-linearity) layout featuring the memory and buffer modes only. b) SQUID imbalance directly appears in the effective circuit found by Foster synthesis of the environment seen by the SQUID. c) Series inductance in the arms of the SQUID enable a two-Cooper pair tunneling mechanism [53] (double-cross junctions) which is not suppressed at the circuit bias point. d) A superinductor made of finite-length chains of Josephson junctions may be viewed as an inductor in parallel of a chain of inductive elements with quartic potential (spider symbols following the convention of Ref. [48]). In general, the effective modes may present a different impedance $\tilde{Z}_{a,b}$ to this element from the imepedance presented to the SQUID. However, given the circuit symmetry, we find $\tilde{Z}_{a,b}=Z_{a,b}$. e) Asymmetric pump waves yield non-linear contributions to the circuit Hamiltonian in the RWA.
@@
We consider the effective two-mode circuit pictured in Fig. S4b, which is obtained after Foster decomposition of the environment seen by the SQUID in the physical circuit (Fig. S4a), and neglecting all modes but the memory and buffer. We derive the ATS inductive potential when the junctions forming its SQUID are imbalanced. Following the convention of Fig. S4b, we let $\varphi$ the reduced phase across the modes in series with the ATS, $\Phi_{1}$, $\Phi_{2}$ the fluxes biasing each loop, $E_{J_{1}}$, $E_{J_{2}}$ the energies of each junction. Defining
@@
($\varphi_{0}=\hbar/(2e)$ is the reduced flux quantum) and using Kirchhoff’s law, we find the ATS potential to be
$V^{ATS}_{\varphi_{\Sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\varphi-\varphi_{\Delta})$ (S8)
@@
where $i=a,b$ (${\bf a}_{i}$ denoting the annihilation operator of mode $i$), $\bm{\varphi}=\sum_{i}\varphi^{ZPF}_{i}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ and $\varphi^{ZPF}_{i}=\big{(}\pi Z_{i}/R_{q}\big{)}^{1/2}$. The cosine term in this Hamiltonian features non-linear, photon-number preserving terms that survive the RWA. At leading order (truncating the cosine at fourth order), we find that each mode inherits a self-Kerr non-linearity of amplitude $\chi_{aa}=\frac{1}{2}E_{\Delta}\varphi_{a}^{ZPF~{}4}$, $\chi_{bb}=\frac{1}{2}E_{\Delta}\varphi_{b}^{ZPF~{}4}$ and both modes are dispersively coupled through a cross-Kerr term $\chi_{ab}=E_{\Delta}\varphi_{a}^{ZPF~{}2}\varphi_{b}^{ZPF~{}2}$. Given the near perfect symmetry of the ATS in our experiment, these terms are negligible ($\chi_{aa},\chi_{bb},\chi_{ab}<2\pi\times$10 kHz).
@@
The dipole energy reads $V=-E_{J}\cos(\psi_{J})+\frac{\epsilon_{L}}{2}\psi_{L}^{2}$. Given the hierarchy $E_{J}\gg\epsilon_{L}$, we can solve the equation (S10) to leading order in $E_{J}/\epsilon_{L}$ and inject the solution in the dipole potential to get
@@
When replacing standard junctions with two such dipoles in the circuit ATS (see Fig. S4c), the inductive potential is renormalized as
$\tilde{V}^{ATS}_{\varphi_{\sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\varphi-2\varphi_{\Delta})$ (S12)
@@
We fit the memory energy spectrum with this model (including corrections from the Josephson junctions asymmetry and a non-linear contribution from the circuit superinductors, see following section), $e_{\Sigma}$ being the only free fit parameter. We thus estimate $e_{\Sigma}$ and finally $\epsilon_{L}=h\times 1920~{}$GHz. This inductive energy corresponds to a series inductance of 85 pH in each arm of the SQUID, which matches a rough estimate of the geometric inductance of the tracks (each arm is 120 $\mu$m long, see Fig. S1). We note that a recent paper [55] reported deviations from the standard $\cos(\varphi)$ potential for junctions embedded in several transmon circuits, with significant higher-order harmonics. This effect could contribute to the $\cos(2\varphi)$ term observed in our experiment. We can nevertheless place a conservative upper bound at 0.2 % for the ratio of the second to first harmonic in the junction potential (corresponding to the case where the $\cos(2\varphi)$ term is entirely attributed to intrinsic distortion of the junction potential). This bound is an order of magnitude smaller than the values reported in [55]. We note that our Josephson junctions are fabricated with a long, high-pressure oxidation step (see Sec. S1.2), probably resulting in an oxide barrier thicker than for the devices analyzed in that work. We can thus conclude, in line with the model presented in [55], that junctions with a thick oxide barrier are less prone the formation of pinholes and are described by a nearly pure $\cos(\varphi)$ potential.
@@
For the data presented in Fig. 3c, reduction of the system Hilbert space is less obvious. Here, we are interested in the steady-state value of $\langle{\bf b}\rangle$. We directly apply the Lindblad adjoint formalism to encode the evolution $\langle{\bf b}\rangle$ along with the expectation value of ten other orthogonal operators (which we compute numerically and do not detail here) in a $11\times 1$ 11 matrix $M$. The steady-state is directly found from the single vector in the kernel of $M$.
@@
We attribute this effect to hybridization of the system to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [57, 58]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, the probability that at least one of the memory or buffer transitions be affected by the TLS bath was high enough that waiting for the bath to reconfigure in a more favorable configuration (either passively or cycling the system to room temperature) was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the circuit transitions away from any resonant coupling with TLS’s.

### v2

@@ V Conclusion and outlook
We demonstrate the engineering of a four-photon dissipation channel whose rate far exceeds the intrinsic single-photon loss rate of the target oscillator. Our approach relies on a near Kerr-free Josephson mixing element embedded in a high-impedance environment, mediating genuine high-order interactions while inducing only residual non-linearity on the involved modes. Even though this non-linearity is larger than the engineered dissipation rate in our experiment—preventing us from dissipating quartets of photons from more than a specific oscillator state—we show that it dominantly originates from a renormalization of the SQUID inductive potential by stray inductances in its tracks. We thus estimate that simple design changes would decrease the circuit non-linearity down to a level where it can be overcome by the four-photon dissipation, enabling the dissipative stabilization of a four-legged cat manifold.
@@ Acknowledgments
The authors thank S. Jezouin, R. Lescanne, A. Murani, and W.C. Smith for fruitful discussions on circuit design, J. Palomo, A. Pierret, M. Rosticher and J.-L. Smirr for assistance in microfabrication, and M. Mirrahimi, P. Rouchon and L.-A. Sellem for valuable discussions on dissipation engineering. Our circuit was fabricated within the consortium Salle Blanche Paris Centre and at the Collège de France. This work has been funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (project Dancingfool, grant agreement No. 101042304 and project Q-feedback grant agreement No. 884762), by the French ANR-22-PETQ-0003 grant and by the ANR-22-PETQ-0006 grant under the ‘France 2030’ plan.
@@
with $\gamma_{4}=4g_{4}^{2}/\kappa_{b}$, and the coherent states spanning the cat manifold are attractors of the dynamics. Note that this property remains true in the regime $g_{4}\gtrsim\kappa_{b}$, even though the two-mode dynamics is no longer accurately captured by (S2[31]. In Eq.(S1), the term $\alpha^{4}{\bf b^{\dagger}}+\mathrm{h.c.}$ is trivially implemented with a resonant drive on the buffer, so that we now focus on the engineering of the more challenging four-to-one photon exchange interaction.
In practice, we engineer this interaction using a pumped ATS participating in the memory and buffer (see Fig. 1). After Foster decomposition of the linear circuit connected to the ATS [41, 25], we reach the Hamiltonian:
@@
If the system is pumped at the frequency $\epsilon_{p}(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$ and after selecting resonant terms assuming $\omega_{a}$, $\omega_{b}$ incommensurate, we get in the RWA the static Hamiltonian
$\tilde{{\bf H}}_{4}={\hbar}g_{4}:P_{a}({\bf a^{\dagger}}{\bf a})P_{b}({\bf b^{\dagger}}{\bf b})~{}{\bf a}^{4}{\bf b^{\dagger}}:~{}+~{}\mathrm{h.c.}$ (S5)
@@
$\begin{split}P_{a}(X)={4!}\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{a}^{ZPF}}^{2n}}{n!(n+4)!}X^{n}\\ P_{b}(X)=\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{b}^{ZPF}}^{2n}}{n!(n+1)!}X^{n}\end{split}$ (S6)
@@
We here note that the energy of each junction forming the SQUID is 3 times larger than its design value (see Table 1), indicating that something unexpected occurred in fabrication. While we cannot identify for certain the origin of the problem, the fact that the memory and buffer mode frequencies are off by only 10 % (indicating that the superinductances have a value 20 % higher than intended) whereas the inductance of the SQUID junctions are lower than intended by 60 % points towards a mechanism that affects more strongly junctions of small area than larger ones. We thus suspect that the electron-beam was poorly focused during patterning.
@@
Figure S4: Sources of non-linearity. a) Ideal circuit (no source of non-linearity) layout featuring the memory and buffer modes only. b) SQUID imbalance directly appears in the effective circuit found by Foster synthesis of the environment seen by the SQUID. c) Series inductance in the arms of the SQUID enable a two-Cooper pair tunneling mechanism [47] (double-cross junctions) which is not suppressed at the circuit bias point. d) A superinductor made of finite-length chains of Josephson junctions may be viewed as an inductor in parallel of a chain of inductive elements with quartic potential (spider symbols following the convention of Ref. [41]). In general, the effective modes may present a different impedance $\tilde{Z}_{a,b}$ to this element from the impedance presented to the SQUID. However, given the circuit symmetry, we find $\tilde{Z}_{a,b}=Z_{a,b}$. e) Asymmetric pump waves yield non-linear contributions to the circuit Hamiltonian in the RWA.
@@
We consider the effective two-mode circuit pictured in Fig. S4b, which is obtained after Foster decomposition of the environment seen by the SQUID in the physical circuit (Fig. S4a), and neglecting all modes but the memory and buffer. We derive the SQUID inductive potential when the junctions forming its SQUID are imbalanced. Following the convention of Fig. S4b, we let $\varphi$ the reduced phase across the modes in series with the SQUID, $\Phi_{1}$, $\Phi_{2}$ the fluxes biasing each loop, $E_{J_{1}}$, $E_{J_{2}}$ the energies of each junction. Defining
@@
($\varphi_{0}=\hbar/(2e)$ is the reduced flux quantum) and using Kirchhoff’s law, we find the SQUID potential to be
$V^{{SQUID}}_{\varphi_{\Sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\varphi-\varphi_{\Delta})$ (S8)
@@
where $i=a,b$ (${\bf a}_{i}$ denoting the annihilation operator of mode $i$), $\bm{\varphi}=\sum_{i}\varphi^{ZPF}_{i}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ and $\varphi^{ZPF}_{i}=\big{(}\pi Z_{i}/R_{q}\big{)}^{1/2}$. The cosine term in this Hamiltonian features non-linear, photon-number preserving terms that survive the RWA. At leading order (truncating the cosine at fourth order), we find that each mode inherits a self-Kerr non-linearity of amplitude $\chi_{aa}=\frac{1}{2}E_{\Delta}{\varphi_{a}^{ZPF}}^{4}$, $\chi_{bb}=\frac{1}{2}E_{\Delta}{\varphi_{b}^{ZPF}}^{4}$ and both modes are dispersively coupled through a cross-Kerr term $\chi_{ab}=E_{\Delta}{\varphi_{a}^{ZPF}}^{2}{\varphi_{b}^{ZPF}}^{2}$. Given the near perfect symmetry of the ATS in our experiment, these terms are negligible ($\chi_{aa},\chi_{bb},\chi_{ab}<2\pi\times$10 kHz).
@@
The dipole energy reads $V=-E_{J}\cos(\psi_{J})+\frac{\epsilon_{L}}{2}\psi_{L}^{2}$. Given the hierarchy $E_{J}\ll\epsilon_{L}$, we can solve the equation (S10) to leading order in $E_{J}/\epsilon_{L}$ and inject the solution in the dipole potential to get
@@
When replacing standard junctions with two such dipoles in the circuit SQUID (see Fig. S4c), the inductive potential is renormalized as
$\tilde{V}^{{SQUID}}_{\varphi_{\sigma},\varphi_{\Delta}}(\varphi)=-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\varphi-\varphi_{\Delta})+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\varphi-2\varphi_{\Delta})$ (S12)
@@
We fit the memory energy spectrum with this model (including corrections from the Josephson junctions asymmetry and a non-linear contribution from the circuit superinductors, see following section), $e_{\Sigma}$ being the only free fit parameter. We thus estimate $e_{\Sigma}$ and finally $\epsilon_{L}=h\times 1920~{}$GHz. This inductive energy corresponds to a series inductance of 85 pH in each arm of the SQUID, which matches a rough estimate of the geometric inductance of the tracks (each arm is 120 $\mu$m long, see Fig. S1). We note that a recent paper [69] reported deviations from the standard $\cos(\varphi)$ potential for junctions embedded in several transmon circuits, with significant higher-order harmonics. This effect could contribute to the $\cos(2\varphi)$ term observed in our experiment. We can nevertheless place a conservative upper bound at 0.2 % for the ratio of the second to first harmonic in the junction potential (corresponding to the case where the $\cos(2\varphi)$ term is entirely attributed to intrinsic distortion of the junction potential). This bound is an order of magnitude smaller than the values reported in [69]. We note that our Josephson junctions are fabricated with a long, high-pressure oxidation step (see Sec. S1.2), probably resulting in an oxide barrier thicker than for the devices analyzed in that work. We can thus conclude, in line with the model presented in [69], that junctions with a thick oxide barrier are less prone to the formation of pinholes and are described by a nearly pure $\cos(\varphi)$ potential.
@@
For the data presented in Fig. 3c, reduction of the system Hilbert space is less obvious. Here, we are interested in the steady-state value of $\langle{\bf b}\rangle$. We directly apply the Lindblad adjoint formalism to encode the evolution of $\langle{\bf b}\rangle$ along with the expectation value of ten other orthogonal operators (which we compute numerically and do not detail here) in a $11\times 1$ 11 matrix $M$. The steady-state is directly found from the single vector in the kernel of $M$.
@@
We attribute this effect to hybridization of the system to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [49, 50]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, the probability that at least one of the memory or buffer transitions be affected by the TLS bath was high enough that waiting for the bath to reconfigure in a more favorable configuration was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the circuit transitions away from any resonant coupling with TLS’s.